OpenStudy (anonymous):

A tower in Maintown sends radio signal a certain distance (in miles) according to the equation: x2 + y2 = 8,100. The Smith family lives 50 miles south and 70 miles east of Maintown. Can the Smith family receive the radio signal from Maintown? In complete sentences, explain your reasoning.

OpenStudy (anonymous):

70^2 + 50^2 = 4,900 + 2,500 <= 8100. So yes.